A set of forces are acting on an object. The first force, with a magnitude of 24.70 N, is directed at ─258o and

the second force, with a magnitude of 53.20 N, is directed at 37o north of east. The third force with a magnitude

of 166.50 N is directed at 270o, and the fourth force of 49.92 N is directed at 50o east of south. Solve for the

resultant force and its exact direction analytically.


1
Expert's answer
2021-02-25T17:46:23-0500

Answer

Components of all force in x direction is written as

"F_x=-24.70sin12^\u00b0+53.20cos37^\u00b0+0+49.92sin50^\u00b0"


"F_x=" -5. 14+42.49+38.24=75.59N

Now y component

"F_y=24.70cos12^\u00b0+53.20sin37^\u00b0-166.50-49.92cos50^\u00b0"

"F_y" =24.16+32.02-166.5-32.09

=-142.41N

Now resultant force

"F=\sqrt{F_x^2+F_y^2}"

="=\sqrt{(75.59)^2+(-142.41)^2}"

=161N

Direction

"\theta=\tan^{-1}(\frac{F_y}{F_x})"

"=\tan^{-1}(\frac{-142.41}{75.59})"

"=-62.11^\u00b0"